OpenStudy (anonymous):

Find the average value of f(t)=−2t^(e−(t^2)) on the interval [0,7]

hartnn (hartnn):

same formula

OpenStudy (anonymous):

i guess −2t^(e(−t^2))

OpenStudy (anonymous):

ok thanks haha

hartnn (hartnn):

then you'l need to substitute u = -t^2

OpenStudy (anonymous):

\[\frac{1}{7 - 0}\int\limits_{0}^{7} f(t)\:dt\]

OpenStudy (anonymous):

thanks guys